OpenStudy (dls):

A machine gun is mounted on the top of tower 100m high.At what angle should the gun be inclined to cover a max range of firing the ground below?Muzzle speed=150m/s,take g=10m/s^2

OpenStudy (dls):

Time is 4.51 s from what i've got if im correct

OpenStudy (dls):

Working with rough approximations..if theta=45,Rmax=2250m,it should be nearby thatonly..iguess

OpenStudy (dls):

@experimentX

OpenStudy (experimentx):

let it be inclined at an angle \( \theta \), the flight time is calculate the total flight time.

OpenStudy (experimentx):

\[ - 2 v \sin \theta/10 + 100/(v \sin \theta)\]

OpenStudy (experimentx):

the horizontal range is given by \[ f(\theta) = v \cos \theta * (- 2 v \sin \theta/10 + 100/(v \sin \theta)) \] maximize this function for theta

OpenStudy (experimentx):

* ... -2v is +2v sin ..

OpenStudy (anonymous):

Corrcect answer (is different) Denote \[v = v_0*Sin \theta\] Then flight time is (assuming that g = 10)\[(-v + \sqrt{20H + v^2})/10 + 2v/10 = 0.1(v + \sqrt{20H + v^2})\] 2v/10

OpenStudy (anonymous):

Now during this time the projectile will cover the distance \[V_0*Cos \theta * 0.1(v+\sqrt{20H + v^2}) = D(\theta,V_0)\]

OpenStudy (anonymous):

So You will have to differentiate with respect to theta the following expression \[V_0 Cos \theta* \left( V_0 Sin \theta + \sqrt{H^2 + V_0^2 Sin^2 \theta}\right)\]

OpenStudy (dls):

Did this question require so many efforts lol :/

OpenStudy (anonymous):

The solution I provided is the most straightforward - start differentiating it

OpenStudy (dls):

im not taught that much high standard atm

OpenStudy (experimentx):

seems 43.664 is the ans

OpenStudy (anonymous):

No "king's road" to correct answer - either sweat or nothing

OpenStudy (anonymous):

Wolfram will A) will do him NO GODD on exam, B) will leave him WITHOUT KNOWLEDGE HOW TO SIMILAR PROBLEMS IN THE FUTURE

OpenStudy (experimentx):

this is just calculus ... maximizing and minimizing. let me see if i can find any other method.

OpenStudy (anonymous):

NO GOOD

OpenStudy (anonymous):

Lets see indeed ..... .... VERY interesting indeed ..................

OpenStudy (dls):

x=u^2sin2theta/g +100cot theta

OpenStudy (anonymous):

Here, u = 150 ms-1 And, Taking g = 10 m-2 (approximately) Let 'θ' be the angle at which the machine gun should fire in order to cover a maximum distance. Then , the horizontal component of velocity = 150 cos θ And, the vertical component of velocity = 150 sin θ If 'T' is the time of flight, then , Horizontal range, R = (150 cos θ ) x T . The gun is mounted at the top of a tower 100 meters high . Let us regard the positive direction of the position-axis as to be along the line from the top of tower in downward direction. For motion along vertical : Initial Velocity = -150 sin θ; Distance covered = + 100 m And, acceleration = + 10 ms-2 In time 'T' , the machine gun shot will reach maximum height and then reach the ground. Now, S = ut + 1/2 at2 Therefore, +100 = ( -150 sin θ ) T + 1/2 x T2 Or, T2 - ( 30 sin θ )T - 20 = 0 Therefore, T = - ( - 30 sin θ ) ± { ( - 30 sin θ)2 - 4 x 1 x (-20) }1/2/2 = 30 sin θ ± (900 sin2 θ + 80 )1/2/2 Or, T = 15 sinθ ± (225 sin2θ + 20)1/2 Now, range will be maximum, if time of flight is maximum . Therefore, choosing positive sign ,we have T = 15 sin θ + ( 225 sin2θ+ 20 )1/2 Hence, horizontal range covered, R = 150 cos θ {15 sin θ+ ( 225 sin2θ + 20 )}1/2 The horizontal range is maximum, when θ = 45o. But in the present case , the machine gun is mounted at height of 100 m . Therefore, R will not be maximum for θ = 45o.It will be maximum for some value of θ close to 45o. If we calculate values of R by setting θ = 43o, 43.5o, 44o, 45o, 46o and 47o, the values of R come out to be 2347 m, 2347.7 m, 2348 m, 2346 m, 2341 m and 2334 m respectively. Thus R is maximum for value of θ some where between 43.5o and 44o. Therefore , the mean value of θ = (43.5o + 44o)/2 = 43.75o. The gun should be inclined at 43.75o to cover a maximum range of firing on the ground below Read more: http://wiki.answers.com/Q/A_machine_gun_is_mounted_on_the_top_of_tower_100m_high_at_what_angle_should_the_gun_be_inclined_to_cover_a_max._range_of_firing_on_ground_below_The_muzzle_speed_of_bullet_is_150_meters_p_sec#ixzz26X9WBUbQ

OpenStudy (dls):

IReadThat

OpenStudy (experimentx):

\[ f(\theta) = v \cos \theta * (+2 v \sin \theta/10 + 100/(v \sin \theta)) \\ v^2 \sin 2\theta /10 + 100 \cot \theta \] doesn't seem easy ... probably you have to resort to calculus.

OpenStudy (anonymous):

THIS STATEMENT IS NOT TRUE "Now, range will be maximum, if time of flight is maximum ."

OpenStudy (anonymous):

@experimentX AS i have said above

OpenStudy (dls):

guys Ive..got...a...simple..solution..here...just..exaplain..me!

OpenStudy (anonymous):

I wanted to write g - but @experimentX here spoiled us by writing 10 everywhere, so I followed his bad example. And in differentiating I just dropped 0.1 (1/g) because it does not change the solution

OpenStudy (dls):